OpenStudy (anonymous):

Devaughn's age is two times Sydney's age. The sum of their ages is 51 . What is Sydney's age?

OpenStudy (anonymous):

let D stand for devaughn and let S stand for sydney so it implies that as at now since D is twice the age of S, we have an equation like D=2S....eqn 1 since their age sum to 51 we have an equation like S+D = 51......eqn2 but we previously have D=2S so we get equation 2 to become S+2S=51 so 3S=51 and therefore S = 51/3 and D = 2times 51/3 which is 102/3...hope it helps